# A bound on the quantum value of all compiled nonlocal games

### v1

@@ 1.2. Comparison with prior work
It is instructive to compare our result with prior works that established quantum soundness in special cases. As mentioned earlier, [NZ23] showed that the compiled value of the CHSH game was bounded by the quantum value. This result was extended to all XOR nonlocal games in [CMM+24, BVB+24]. The fact that these results appear to give tighter bounds on the compiled value by the quantum value $\omega_{q}$ rather than the commuting operator value $\omega_{qc}$, but in fact both values coincide for the XOR case by a result of Tsirelson [Tsi87]. This coincidence is also apparent in the SOS proof techniques in [NZ23, CMM+24, BVB+24], which naturally relate to the commuting operator value rather than the quantum one. Nonlocal games with a sequential (temporal) separation rather than a spatial one have also been investigated in the XOR case [CFE+24].
1.3. Organization of the paper
@@
The remainder of the paper is organized as follows: Section 2 covers preliminary material. Section 3 provides an overview of nonlocal games, correlations, and the various values of those games in the spatially separated setting. Section 4 details the KLVY compiler and the description of a compiled nonlocal game, including a discussion of the value of a compiled nonlocal game, and proves a key technical result. Section 5 establishes our equivalent characterization of commuting operator correlations (as well as of classical and quantum correlations) in terms of strategies for a sequential game that satisfy a strong non-signaling property. Section 6 proves of our main result, establishing the upper bound on the quantum value of a compiled nonlocal game.
2. Preliminaries
@@
2.1. Vectors, operators, quantum mechanics
@@
In quantum mechanics, physical systems are often identified with Hilbert spaces $\mathcal{H}$, and the states of the system are identified with positive semidefinite operators $\rho$ with unit trace, called density operators. A state is called pure if the density operator has rank one, and otherwise it is called mixed. Any unit vector $\lvert v\rangle\in\mathcal{H}$ determines a pure state by the formula $\rho=\lvert v\rangle\langle\mathopen{}v\rvert$, and conversely any pure state can be written in this way, hence the two concepts are often identified. The trace distance is the statistical distance between the distributions associated with two density operators $\rho$ and $\sigma$ is given by the formula $\frac{1}{2}\lVert\rho-\sigma\rVert_{1}=\frac{1}{2}\operatorname{tr}(\lvert\rho-\sigma\rvert)$, where $\lVert\cdot\rVert_{1}$ is the Schatten-$1$ norm and the absolute value of an operator is defined by $\lvert A\rvert:=\sqrt{A^{*}A}$. A measurement with a finite outcome set $\mathcal{O}$ is described by a collection of bounded operators $\{A_{a}\}_{a\in\mathcal{O}}$ acting on $\mathcal{H}$ such that $\sum_{a\in\mathcal{O}}A_{a}^{*}A_{a}=\mathbbm{1}$. If the system is in state $\rho$, then the probability of obtaining outcome $a$ is given by $p(a)=\operatorname{tr}(A_{a}^{*}A_{a}\rho)$, after which the state of the system is described by $A_{a}\rho A_{a}^{*}/p(a)$. The probabilities of measurement outcomes only depend on the operators $M_{a}:=A_{a}^{*}A_{a}$. A collection of operators $\{M_{a}\}_{a\in\mathcal{O}}$ such as these which satisfy $\sum_{a\in\mathcal{O}}M_{a}=\mathbbm{1}$ is called a POVM, which is short for positive operator-valued measure, with outcomes in $\mathcal{O}$. Any POVM arises from a measurement. Observables are self-adjoint elements $B=B^{*}\in\mathcal{B}(H)$, and their quantum expectation value with respect to the state $\rho$ is given by $\operatorname{tr}(\rho B)$. This can be related to the preceding if one takes $\mathcal{O}$ to be the set of eigenvalues of $B$ (assuming it is finite) and $A_{a}$ as the corresponding spectral projections. We will often discuss apparatuses with multiple measurement settings, labeled by some index set $\mathcal{I}$, but the same set of outcomes $\mathcal{O}$ for each setting. This will be denoted by $\{\{M_{ax}\}_{a\in\mathcal{O}}:x\in\mathcal{I}\}$, where $\{M_{xa}\}_{a\in\mathcal{O}}$ is be a POVM (or measurement) with outcomes in $\mathcal{O}$ for each $x\in\mathcal{I}$. We often abbreviate and write this as $\{M_{xa}\}_{a\in\mathcal{O},x\in\mathcal{I}}$ when clear from context.
2.2. Algebras and representations

### v3

@@ 1.2 Main results
It is instructive to compare our general bound with prior works on compiled nonlocal games that established quantum soundness in special cases – first for the CHSH game and then for general XOR games [33, 15, 2]. Although these results appear to give tighter bounds on the compiled value, bounding it by the quantum value $\omega_{q}$ rather than the commuting operator value $\omega_{qc}$, this is not in fact the case – by a result of Tsirelson [44], we have $\omega_{q}=\omega_{qc}$ for the class of XOR games! This coincidence is also apparent in the SOS proof techniques in [33, 15, 2], which naturally bound the commuting operator value rather than the quantum one, but so far have resisted generalization to general games.
Self-testing is a powerful primitive in many applications of nonlocal games, as it allows inferring the prover’s strategy from the observed correlations (up to unobservable degrees of freedom). It has both theoretical and practical significance. For example, a self-testing result for the compiled CHSH game was key ingredient for BQP verification [33].
While self-testing is conventionally defined in terms of local dilations [27], this is not suitable for the commuting operator setting. Following [38], we say a nonlocal game $\mathcal{G}$ is a commuting operator self-test if any optimal commuting operator strategy gives rise to the same expectation values, for all polynomials in the POVM elements. The latter can be conveniently captured in terms of a state on the max-tensor product of abstract POVM algebras (Definition 6.3). We can then prove the following asymptotic self-testing result in the compiled setting:
Theorem 1.2.
Let $\mathcal{G}$ be a commuting operator self-test. If a QPT prover wins the compiled game with probability tending to the commuting operator value, then for any question-answer pair for Alice the expectation values of all polynomials in Bob’s POVMs are uniquely determined as $\lambda\to\infty$.
In other words, we show that if $\lim_{\lambda\to\infty}\omega_{\lambda}(\mathcal{G}_{\textnormal{comp}},S)=\omega_{qc}(\mathcal{G})$ then
$\lim_{\lambda\to\infty}\operatorname{tr}\left(\sigma_{xa}^{\lambda}\,P(\{B^{\lambda}_{yb}\})\right)$
is uniquely determined by the optimal commuting operator strategy, for every $x\in\mathcal{I}_{A},a\in\mathcal{O}_{A}$ and for every noncommutative polynomial $P$. Here, $\sigma_{xa}^{\lambda}$ denotes the state of the QPT strategy after the interactions corresponding to question $x$ and answer $a$, and $B_{yb}$ are the POVM elements for question $y$ and answer $b$ applied subsequently (cf. Fig. 1). The proof of Theorem 1.2 follows from the second statement in Theorem 6.5.
While the preceding general discussion is necessarily somewhat abstract, its consequences are concrete. For example, the CHSH game is a commuting operator self-test and it is known that in any optimal strategy the Bob observables anti-commute on the state. Thus our theorem shows that in any asymptotical optimal strategy, it holds that $\lim_{\lambda\to\infty}\operatorname{tr}\left(\sigma_{xa}^{\lambda}\,\{B^{\lambda}_{0},B^{\lambda}_{1}\}^{2}\right)=0$, where $B^{\lambda}_{x}:=B^{\lambda}_{x0}-B^{\lambda}_{x1}$. In this way we recover as a special case a version of the self-testing result established and used to great effect in [33].
1.3 Technical outline
To set some context for our results, let us first recall the basics of the KLVY compiler [24]. Given a two-player nonlocal game $\mathcal{G}$ the KLVY transformation turns it into a single-player compiled game $\mathcal{G}_{\textnormal{comp}}$ as follows:
  • The referee sends to the player an encryption of what would be Alice’s question $\xi=\mathsf{Enc}(x)$, under a QHE scheme.
  • The prover homomorphically computes an answer and returns the corresponding ciphertext $\alpha=\mathsf{Enc}(a)$ to the referee. (In the honest case, this is Alice’s encrypted answer)
  • The referee sends Bob’s question $y$ in the plain.
  • The player returns an answer $b$. (In the honest case, this is Bob’s answer)
It is clear that the quantum value of the compiled game is at least the quantum value of $\mathcal{G}$, since the player can run the optimal strategy for $\mathcal{G}$, by the homomorphic properties of QHE. The more challenging aspect is to prove an upper bound on the quantum value of $\mathcal{G}_{\textnormal{comp}}$, which is the focus of this work.
Let us start with a simple observation about the compiled game. Since Bob’s question is only revealed after Alice’s question was asked and answered, the marginal distribution of Alice’s answer must be independent of Bob’s question. In other words, any player’s strategy must always be non-signaling from Bob to Alice.
For the other direction, we would ideally like to say that the same holds, because of the security of the encryption scheme, which leaks no information about Alice’s question to the player. Unfortunately, this is not quite true and we rather have a weaker form of computational non-signaling, i.e., the state of the player is indistinguishable from a (possibly different) state that is truly independent of Alice’s question. To make this more precise, let $\lambda$ be the security parameter, $\sigma_{x}^{\lambda},\sigma_{x^{\prime}}^{\lambda}$ be the states after Alice’s computation for question $x$ and $x^{\prime}$, respectively, and let $\{B_{yb}^{\lambda}\}$ be the POVM element corresponding to the player’s behavior in the last two rounds (Bob’s computation) for question $y$ and answer $b$. Then computational non-signaling says that:
$\operatorname{tr}(\sigma_{x}^{\lambda}B_{yb}^{\lambda})\approx\operatorname{tr}(\sigma_{x^{\prime}}^{\lambda}B_{yb}^{\lambda})$
up to negligible factors. In fact, we can push this observation one step further, by noting that the parallel/sequential combination of efficient algorithms is once again efficient, and thus the above guarantee should also hold for such algorithms. In more details, block encodings of efficient POVMs behave well under linear combination and multiplication allowing us to make a stronger statement about computational non-signaling. We can show that for all (constant-degree) polynomials $P=P(\{B_{yb}\})$ of non-commuting variables $\{B_{yb}\}$, we have:
$\operatorname{tr}(\sigma_{x}^{\lambda}P(\{B_{yb}^{\lambda}\}))\approx\operatorname{tr}(\sigma_{x^{\prime}}^{\lambda}P(\{B_{yb}^{\lambda}\})).$
We refer to this property as computationally strong non-signaling. By generalizing results from [33, 15] from observables to polynomials of POVMs, we can prove that the security of the QHE scheme guarantees that the strategies in the compiled game are computationally strong non-signaling. The formal result is stated in Proposition 4.6.
To make sense of why this generalized property is useful, let us first consider an idealized scenario without encryption (and hence no security parameter $\lambda$), where we have an exact equality $\operatorname{tr}(\sigma_{x}P(\{B_{yb}\}))=\operatorname{tr}(\sigma_{x^{\prime}}P(\{B_{yb}\}))$. In other words, we consider an (unrealistic and) idealized situation of a sequential game (cf. Fig. 1) that is exactly strongly non-signaling, i.e., where the player’s state after the first two rounds is independent of $x$. Dealing with the additional negligible factor will be an important part of the challenge, but we will come back to it at a later point in this overview.
We claim that any strategy for the sequential game that is guaranteed to be (exact) strongly non-signaling, is in fact a commuting operator strategy in the corresponding nonlocal game. We show a sketch of this proof, for the simplified case of finite-dimensional strategies. This will not be sufficient for our main result (for reasons that will become clear later), but it is useful to gain some intuition on why this statement is actually true.
By definition of strong non-signaling, we have that the (mixed) states $\sigma_{x}$ after the first two rounds of interaction are independent of $x$, meaning that those states are all the same. Let $\sigma:=\sigma_{x}$ for all $x$ denote this state. Let further $\sigma_{xa}$ be the (subnormalized) state acting on a finite-dimensional Hilbert space $\mathcal{H}$ after the first two rounds of interaction, where $x$ is Alice’s question and $a$ is Alice’s answer. Consider an arbitrary purification $\lvert\psi_{xa}\rangle$ of this state, we consequently obtain that:
$\lvert\psi_{x}\rangle=\sum_{a\in\mathcal{O}_{A}}\lvert a\rangle\otimes\lvert\psi_{xa}\rangle$
where $\mathcal{O}_{A}$ is the set of possible Alice answers. Since these states are all purifications of the same state $\sigma$, by Uhlmann’s theorem, they are all related by a unitary operation acting only on Alice’s register. We can use this fact to construct a quantum correlation as follows: Let $\lvert\psi_{x_{0}}\rangle$ as above, for some fixed $x_{0}\in\mathcal{O}_{A}$, be the joint state shared between Alice and Bob. While the Bob operator stays the same as in the sequential game, Alice first applies a unitary $U_{x}$ to switch to the purification $\lvert\psi_{x}\rangle$ and then applies the projection $P_{a}=\lvert a\rangle\langle a\rvert$. This gives us a quantum correlation:
$p(a,b|x,y)=\operatorname{tr}(\sigma_{xa}B_{yb})=\langle\psi_{x_{0}}\rvert U_{x}P_{a}U_{x}^{*}\otimes B_{yb}\lvert\psi_{x_{0}}\rangle$
which is identical to the correlation in the sequential game. The formal result is stated in Propositions 5.12 and 5.11.
As hinted above, this proof is not sufficient for us, precisely because the correlations are not exactly non-signaling, but rather computationally non-signaling. One may wishfully hope that, taking the limit of the security parameter to infinity, one gets rid of the extra negligible function, allowing to execute the above strategy. Unfortunately, this does not appear to be the case, since Bob POVM’s do not necessarily generate the whole algebra of bounded operators on a Hilbert space, and therefore the Hilbert space at its limiting point may not even be well-defined. Instead, what we rather get is that $\sigma_{x}$ and $\sigma_{x^{\prime}}$ are indistinguishable for polynomials of Bob POVM elements, i.e., it holds that $\operatorname{tr}(\sigma_{x}P(\{B_{yb}\}))=\operatorname{tr}(\sigma_{x^{\prime}}P(\{B_{yb}\}))$ for all Alice questions $x,x^{\prime}$. What then does hold, is that Bob’s POVMs (and in particular non-commutative polynomials of Bob’s POVMs) generate a subalgebra of the algebra of bounded operators on some Hilbert space. In operator theory language, this is a so-called $C^{*}$-algebra.
Among other things, the framework of $C^{*}$-algebras offers a model of states and observables in quantum mechanics. In our context, this means that every strategy of the compiled game can be captured by $C^{*}$-algebras: The first two rounds by partial states $\phi_{xa}$ on a $C^{*}$-algebra $\mathscr{B}$, and the last two rounds by POVM elements $B_{yb}$ in $\mathscr{B}$. Such an algebraic strategy naturally gives rise to correlations $p(a,b|x,y)=\phi_{xa}(B_{yb})$. The definition of strong non-signaling also extends to the algebraic setting by requiring that the states $\phi_{x}:=\sum_{a\in\mathcal{O}_{A}}\phi_{xa}$ on the $C^{*}$-algebra $\mathscr{B}$ are the same for every Alice question $x$, i.e. $\phi_{x}=\phi_{x^{\prime}}$. Note that, for the algebra of bounded operators on a Hilbert space $\mathscr{B}=\mathcal{B}(\mathcal{H})$ we recover the previous strongly non-signaling definition.
Guided by this intuition, we can infer that the right statement to prove for the sequential game is that strongly non-signaling algebraic strategies characterize the commuting operator correlations. We show this by relying on two fundamental concepts from the theory of operator algebras:
  • (i)
    To realize an abstract $C^{*}$-algebra concretely on a Hilbert space we use the GNS construction.
  • (ii)
    To put together the commuting operator correlation on the Hilbert space we get from applying the GNS construction, we use the Radon-Nikodym theorem for $C^{*}$-algebras which acts like the purification in the finite-dimensional proof we saw in the beginning.
What is left to be done is to move from the idealized setting to the real setting, where the strong non-signaling property holds only up to a negligible (vanishing) summand in the security parameter $\lambda$. In other words, we can only rely on the fact that, for all $x,x^{\prime}$, the states:
$\sigma^{\lambda}_{x}\approx\sigma^{\lambda}_{x^{\prime}}$
are computationally indistinguishable. Note that these are operators on different Hilbert spaces for different $\lambda$, and the same is true for the POVM elements $\{B^{\lambda}_{yb}\}$ that correspond to the last two rounds. To obtain our main result, we have to analyze what happens when one takes the limit as $\lambda\to\infty$, which we achieve by incorporating mathematical tools from operator algebras.
For any QPT strategy, we have a sequence $\{\omega_{\lambda}\}_{\lambda}$ of winning probabilities for the compiled game $\mathcal{G}_{comp}$, indexed by the security parameter. Since the interval $[0,1]$ is (sequentially) compact and it holds $\omega_{\lambda}\in[0,1]$ for all $\lambda$, we obtain that this sequence of winning probabilities has a convergent subsequence. Our goal is to prove that no QPT strategy can win the compiled game with probability exceeding the quantum commuting operator value of the game by any constant. To show this result, we have to bound the limit of any convergent subsequence by the quantum commuting operator value. While we get a convergent subsequence from the compactness of the interval $[0,1]$, this is not enough to obtain our bound: The values $\omega_{\lambda}$ do not keep track of the states and POVM’s that are present in the QPT strategy. Therefore, we have to use a more advanced compactness argument to achieve our goal.
Specifically, we show that the essential part of the prover’s strategies can be captured by a sequence of states on a single universal $C^{*}$-algebra, and use a compactness argument to prove the existence of a limiting state. We want to work with the $C^{*}$-algebra $\mathscr{A}_{\text{POVM}}^{\mathcal{I}_{B},\mathcal{O}_{B}}$ where $\mathcal{I}_{B}$ and $\mathcal{O}_{B}$ are the corresponding finite question and answer sets of Bob. The algebra is generated by elements $\{B_{yb}\}_{y\in\mathcal{I}_{B},b\in\mathcal{O}_{B}}$ satisfying the relations of a POVM for every $y$ (and in particular the set of non-commutative polynomials of Bob’s POVMs is dense in this algebra). Its most important feature is the following universal property: For any collection of POVMs $\{B_{yb}^{\lambda}\}$ acting on some Hilbert space $\mathcal{H}^{\lambda}$, there is a unique *-homomorphism:
$\theta_{\lambda}:\mathscr{A}_{\text{POVM}}^{\mathcal{I}_{B},\mathcal{O}_{B}}\to\mathscr{B}(\mathcal{H}^{\lambda})$
that maps $B_{yb}\mapsto{B_{yb}^{\lambda}}$ for all $y\in\mathcal{I}_{B}$ and $b\in\mathcal{O}_{B}$. Recall that the correlations of a QPT strategy are given by $p_{\lambda}(a,b|x,y)=\operatorname{tr}(\sigma_{xa}^{\lambda}B_{yb}^{\lambda})$ for some subnormalized states $\sigma_{xa}^{\lambda}$ and POVM’s $B_{yb}^{\lambda}$ on some Hilbert space $\mathcal{H}^{\lambda}$. Now, since $\mathscr{A}_{\text{POVM}}^{\mathcal{I}_{B},\mathcal{O}_{B}}$ has the universal property described above, we can define subnormalized states on the algebra by:
$\phi_{xa}^{\lambda}:\mathscr{A}_{\text{POVM}}^{\mathcal{I}_{B},\mathcal{O}_{B}}\to\mathbb{C}\quad\text{s.t.}\quad\phi_{xa}^{\lambda}(\cdot)=\operatorname{tr}(\sigma_{xa}^{\lambda}\theta_{\lambda}(\cdot)).$
Note that all maps $\phi_{xa}^{\lambda}$ have the same domain now, allowing us to use compactness arguments. Since the norm of each such $\phi_{xa}^{\lambda}$ is bounded by $1$, we can apply the Banach-Alaolgu theorem – a central compactness theorem in functional analysis – to show that the set of those $\phi_{xa}^{\lambda}$ is sequentially compact in the weak-* topology. Therefore any sequence $\{\phi_{xa}^{\lambda}\}_{\lambda}$ has a convergent subsequence $\{\phi_{xa}^{\lambda_{k}}\}_{k}$ that converges pointwise to a functional $\phi_{xa}$. It can be verified easily that $\phi_{x}:=\sum_{a\in\mathcal{O}_{A}}\phi_{xa}$ are indeed states of the $C^{*}$-algebra. The limiting states can be then shown to precisely satisfy strong non-signaling, i.e. $\phi_{x}=\phi_{x^{\prime}}$! Since the limiting correlation is strongly non-signaling, we deduce from the previous paragraph that the limiting correlation is a commuting operator correlation. Therefore, the winning probability of the limiting correlation of any subsequence of $\{\omega_{\lambda}\}_{\lambda}$ can at most be the commuting operator value of the nonlocal game. This concludes the proof of our main result.
The line of reasoning in our proof is reminiscent of proofs of completeness for noncommutative optimization hierarchies such as the NPA hierarchy [35]. It is interesting that it also applies to problems in computationally-secure cryptography, and we expect that it can find more uses in this setting.
We conclude this overview with some additional observations on self-testing. Since $\omega_{q}(\mathcal{G})=\omega_{qc}(\mathcal{G})$ holds for many nonlocal games, our result enables self-testing theorems in the compiled setting. It turns out that our characterization of quantum commuting correlations (Theorem 5.15) already implies such results, if we work with an abstract self-testing definition. In the Hilbert space formulation, a self-test is a nonlocal game for which any optimal quantum strategy is the same as a chosen ideal strategy, up to local isometries. The spirit of this definition can be translated to the abstract $C^{*}$-algebra world where a nonlocal game is called a commuting operator self-test if any commuting operator strategy that achieves the quantum commuting value determines the same state on the $C^{*}$-algebra $\mathcal{A}_{\text{POVM}}^{\mathcal{I}_{A},\mathcal{O}_{A}}\otimes_{\max}\mathcal{A}_{\text{POVM}}^{\mathcal{I}_{B},\mathcal{O}_{B}}$ – the (maximal) tensor product of the algebras generated by the Alice and Bob POVMs.
We prove an asymptotic self-testing statement for the compiled nonlocal game using the abstract self-test formulation. Let $\mathcal{G}$ be a nonlocal game that is a commuting operator self-test with corresponding unique state $\Psi$ on $\mathcal{A}_{\text{POVM}}^{\mathcal{I}_{A},\mathcal{O}_{A}}\otimes_{\max}\mathcal{A}_{\text{POVM}}^{\mathcal{I}_{B},\mathcal{O}_{B}}$. Consider a strongly non-signaling algebraic strategy with subnormalized states $\phi_{xa}$ and POVMs $\{\tilde{B}_{yb}\}$ that achieves the quantum commuting value. It holds that for every noncommutative polynomial $P$, we have:
$\phi_{xa}(P(\{\tilde{B}_{yb}\}))=\Psi(A_{xa}\otimes P(\{B_{yb}\})).$
This is proven by combining the GNS construction with the fact that the nonlocal game is a commuting operator self-test. For every QPT strategy in the corresponding compiled game $\mathcal{G}_{\textnormal{comp}}$ with probability tending to the quantum commuting value, we therefore get that:
$\lim_{\lambda\to\infty}\operatorname{tr}\left(\sigma_{xa}^{\lambda}\,P(\{B^{\lambda}_{yb}\})\right)=\Psi(A_{xa}\otimes P(\{B_{yb}\})).$
This shows uniqueness of the corresponding expectation values, since the state $\Psi$ is unique.
1.4 Open problems and outlook
Our main theorem shows that no QPT strategy can win the compiled game $\mathcal{G}_{\textnormal{comp}}$ with probability exceeding $\omega_{qc}(\mathcal{G})$ by any constant. Since we do not know how to compile general commuting operator strategies, there might be a tighter upper bound. Is $\omega_{qc}(\mathcal{G})$ the true answer, or is it instead the quantum value $\omega_{q}(\mathcal{G})$? Since compiled games are fundamentally finite-dimensional, the former might seem implausible. But recall the result by Ozawa [37] that describes $C_{qc}$ as the limit of approximately commuting finite-dimensional quantum strategies as the commutation error approaches zero; cf. [14]. Since QPT strategies for compiled games give rise to finite-dimensional strategies that similarly relax an information-theoretic property (strong non-signaling instead of commutation), this makes it not completely implausible to consider the tantalizing possibility that $\omega_{qc}(\mathcal{G})$ might be the correct answer. On the other hand, while the QHE scheme allows for compiling tensor product strategies, there is no straightforward way to compile approximately commuting strategies for the KLVY approach. If it turns out that there is no way in doing so, the quantum value $\omega_{q}(\mathcal{G})$ might be an upper bound for the winning probability of any QPT strategy for the compiled game.
Our results characterize the limit $\lambda\to\infty$, but can one establish a quantitative bound? For example, does it converge faster than any polynomial in $\lambda$? This is known for compiled XOR games [15] and would be useful for applications. We speculate that proving this for general games would require substantially different techniques. A particular obstruction is the conjectured undecidability of the gapped quantum commuting operator value problem [31].
Classical soundness was shown for nonlocal games with $k\geq 2$ players [24], quantum soundness was shown for CHSH, respectively XOR games, for $k=2$ players [33, 15, 2] and the soundness results in this paper are proven for all nonlocal games with $k=2$ players. We leave it for future work to generalize our soundness results for $k\geq 2$ players. Furthermore, the KLVY compiler employs a QHE scheme and thus relies on strong cryptographic assumptions. The security of the QHE scheme itself only plays a small part of the proof of quantum soundness of the KLVY compiler (Lemma 4.8). Thus, it seems plausible that our analysis will also work for other compilers with weaker cryptographic assumptions.
1.5 Organization of the paper
@@
The remainder of the paper is organized as follows: Section 2 covers preliminary material. Section 3 provides an overview of nonlocal games, correlations, and the various values of those games in the spatially separated setting. Section 4 details the KLVY compiler and the description of a compiled nonlocal game, including a discussion of the value of a compiled nonlocal game, and proves a key technical result. Section 5 establishes our equivalent characterization of commuting operator correlations (as well as of classical and quantum correlations) in terms of strategies for a sequential game that satisfy a strong non-signaling property. Section 6 contains the proof of our main result, which establishes the upper bound on the quantum value of a compiled nonlocal game, and the proof of our self-testing result that is obtained using the same techniques.
2 Preliminaries
@@
2.1 Vectors, operators, quantum mechanics
@@
In quantum mechanics, physical systems are often identified with Hilbert spaces $\mathcal{H}$, and the states of the system are identified with positive semidefinite operators $\rho$ with unit trace, called density operators. A state is called pure if the density operator has rank one, and otherwise it is called mixed. Any unit vector $\lvert v\rangle\in\mathcal{H}$ determines a pure state by the formula $\rho=\lvert v\rangle\langle\mathopen{}v\rvert$, and conversely any pure state can be written in this way, hence the two concepts are often identified. The trace distance is the statistical distance between the distributions associated with two density operators $\rho$ and $\sigma$ and is given by the formula $\frac{1}{2}\lVert\rho-\sigma\rVert_{1}=\frac{1}{2}\operatorname{tr}(\lvert\rho-\sigma\rvert)$, where $\lVert\cdot\rVert_{1}$ is the Schatten-$1$ norm and the absolute value of an operator is defined by $\lvert A\rvert:=\sqrt{A^{*}A}$. A measurement with a finite outcome set $\mathcal{O}$ is described by a collection of bounded operators $\{A_{a}\}_{a\in\mathcal{O}}$ acting on $\mathcal{H}$ such that $\sum_{a\in\mathcal{O}}A_{a}^{*}A_{a}=\mathbbm{1}$. If the system is in state $\rho$, then the probability of obtaining outcome $a$ is given by $p(a)=\operatorname{tr}(A_{a}^{*}A_{a}\rho)$, after which the state of the system is described by $A_{a}\rho A_{a}^{*}/p(a)$. The probabilities of measurement outcomes only depend on the operators $M_{a}:=A_{a}^{*}A_{a}$. A collection of operators $\{M_{a}\}_{a\in\mathcal{O}}$ such as these which satisfy $\sum_{a\in\mathcal{O}}M_{a}=\mathbbm{1}$ is called a POVM, which is short for positive operator-valued measure, with outcomes in $\mathcal{O}$. Any POVM arises from a measurement. Observables are self-adjoint elements $B=B^{*}\in\mathcal{B}(H)$, and their quantum expectation value with respect to the state $\rho$ is given by $\operatorname{tr}(\rho B)$. This can be related to the preceding if one takes $\mathcal{O}$ to be the set of eigenvalues of $B$ (assuming it is finite) and $A_{a}$ as the corresponding spectral projections. We will often discuss apparatuses with multiple measurement settings, labeled by some index set $\mathcal{I}$, but the same set of outcomes $\mathcal{O}$ for each setting. This will be denoted by $\{\{M_{xa}\}_{a\in\mathcal{O}}:x\in\mathcal{I}\}$, where $\{M_{xa}\}_{a\in\mathcal{O}}$ is a POVM (or measurement) with outcomes in $\mathcal{O}$ for each $x\in\mathcal{I}$. We often abbreviate and write this as $\{M_{xa}\}_{a\in\mathcal{O},x\in\mathcal{I}}$ when clear from context.
2.2 Algebras and representations